If a savings account pays interest at 4% compounded quarterly, then the amount of $1 left on deposit for 8 years would be found in a table using:

? periods at ?%.

.
Answer #1

Solution:

Given data:

  • Rate of interest = 4% compounded quarterly
  • Time period = 8 yrs

Hence,interest is 4% compounded quarterly so it will divided by 4 (i.e.number of quarters in a year)

then interest is 4/4 = 1%

Deposit period is 8 yrs , it is multiplied by 4 then total no. of periods = 8*4 =32 periods.

Therefore, the answer is 32 periods at 1%